# CoSQA+: Pioneering the Multi-Choice Code Search Benchmark with Test-Driven Agents

### v4

@@ Introduction
The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matching codes from a large codebase. This forms the foundation for annotating query-code pairs in the CoSQA+ dataset. Firstly, we employ three well-established code embedding models to generate vector representations for both queries and codes. Then, these embeddings are used to calculate cosine similarity scores, averaging similarity results across the three models. Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query. For a detailed breakdown of query distributions by the number of exactly matched codes, please see Appendix D. Distribution of Queries Based on Exactly Matched Codes.
@@
High similarity between query and code is fundamental for a high-quality dataset  (Sun et al., 2022). During our preliminary assessments, we observe disparities in the similarity scores calculated by different models. Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably. And this will potentially lead to unfair test result of benchmark. To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs. Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets. Specifically, we use three models: CodeBERT  (Feng et al., 2020), UniXcoder  (Guo et al., 2022) and CodeT5+ 110M embedding  (Wang et al., 2023).
@@
By taking the average similarity score from these three models, we promote a more balanced and fair assessment of query-code pair quality.
@@
Before constructing the full set of candidate pairs, we randomly selected 2,000 queries and annotated them using test-driven agents to determine the optimal number of candidates per query. The experimental results indicate that the probability of identifying relevant codes decreases significantly as the number of candidates increases (Figure 2). As shown in our analysis, when the number of candidates reaches 20, the marginal benefit of further expansion diminishes. This decline can be attributed to limitations of codebase, nature of queries and ability of embedding models. Furthermore, adding just one more candidate code per query results in an additional 20,604 pairs, which would incur substantial costs in terms of both time and financial resources for annotation. Thus, limiting candidates to the top-20 strikes a balance between maximizing the retrieval of relevant codes and controlling annotation costs.
@@
Test-Driven Agents Annotation
To efficiently and accurately annotate the large number of query-code pairs, we design a test-driven annotation agents based on DeepSeek-V3. An preliminary screener will determine whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment. For uncertain cases, we determine the labels based on the test execution results and final judgment. In both paths, labels are assigned to indicate whether the code matches the query’s requirements (1) or not (0). All labeled query-code pairs are collected to form our dataset CoSQA+, which consists of pairs that have been rigorously verified either through direct assessment or test-based validation. All prompts and a detailed example are provided in AppendixA.
@@
The evaluation process begins with a preliminary screener that categorizes query-code pairs into three distinct labels: (1) clear match, (2) clear non-match, and (3) unclear, as illustrated in Figure 3. Guided by the prompt instructions, the model assesses whether the code fully addresses the query’s functional requirements. For unambiguous cases, it provides a justification and immediate classification. For ambiguous cases (representing 33% of candidates), it flags the pair for deeper verification by downstream agents. This triage mechanism optimizes the annotation workflow by resolving straightforward cases upfront while reserving computational resources for ambiguous scenarios.
@@
For ambiguous cases, we employ a test-driven verification methodology. The test program generator constructs executable test suites designed to rigorously validate code functionality. Each test suite includes dependency scaffolding — importing required libraries, defining auxiliary classes, and declaring helper functions — to ensure runtime compatibility. The tests incorporate assert-based validation mechanisms targeting edge cases, functional requirements, and behavioral constraints specified in the query. Considering that a single query may have multiple valid explanations, the test program is generated based on one of the plausible interpretations of the query. This approach transforms ambiguous cases into executable specifications for objective validation.
@@
The framework ultimately concludes with a final arbiter stage, where the agent employs a specialized prompt to assess if the generated code fully satisfies the query. This assessment is based on both the test program produced and the outcomes of its execution. This methodology shifts verification from theoretical code inspection to empirical observation, empowering the agent to establish evidence-based conclusions by systematically evaluating actual runtime behavior. By grounding judgments in dynamic test execution rather than relying exclusively on static structural analysis and semantic analysis. An example is provided in Appendix BB. An example of Final Arbiter.
Benchmark Statistics
In total, 412,080 labeled query-code pairs are retained. Table 1 shows the statistics of our CoSQA+ benchmark.
Number of pairs Number of code Code average length Average test cases Executable rate
@@
Table 1: Statistics of CoSQA+.
Cost Analysis
@@
Evaluation
@@
RQ1: Quality Comparison between CoSQA+ and CoSQA
@@
Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG  (Wang et al., 2013) and MRR  (Feng et al., 2020). NDCG measures the quality of ranking by considering both the relevance and position of retrieved results. For a given query $q_{i}$, we calculate NDCG@k as:
@@
In the first experiment, we randomly select 1000 query-code pairs labeled "1" from CoSQA+ and CoSQA, ensuring the queries are identical. For CoSQA+, a query may have multiple correct codes, so we randomly select one that differs from the corresponding query’s code in CoSQA. These queries are then presented to experienced programmer volunteers, who are asked to choose the code that satisfies more demand of the query. Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
@@
Table 2: Performance of code search models fine-tuned with different data. The test set is CSN Python.
@@
After the volunteers compare the two datasets, we come to the result that out of the 1000 selected queries, 73.8% of the better codes are from CoSQA+, while 26.2% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table 2, the three models fine-tuned with CoSQA+ outperforms those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in NDCG@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
@@
RQ2: Performance of Agent-based Annotator
@@
To establish a reliable ground truth, we first have three Python experts design test programs for 300 randomly selected query-code pairs and annotate them based on test execution results. The final ground truth labels are determined by majority voting among these three experts. We then evaluate three different annotation approaches against this ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet).
For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators.
@@
Agent-based Annotator 0.964
Human Wo. Test Program 0.890
Deepseek-V3 0.820
O1-mini 0.816
Claude-3.5-Sonnet 0.830
Table 3: Accuracy comparison of different annotation methods against ground truth.
As shown in Table 3, our agent-based annotator achieves 96.4% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (82.0%), O1-mini (81.6%) and Claude-3.5-Sonnet (83.0%). This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification.
@@
In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.85, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
RQ2 Summary: Our agent-based annotator achieves 96.4% accuracy, outperforming both traditional human annotation and single LLM approaches. The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
RQ3: Existing Methods Performance on CoSQA+
@@
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using NDCG@10 as the primary metric. NDCG@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match. Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly.
The tested methods include both traditional information retrieval techniques and advanced deep learning-based models. Specifically, we compare a traditional Bag of Words (BoW) approach (Zhang et al., 2010) with state-of-the-art semantic code search models, namely: CodeBERT (Feng et al., 2020), UniXcoder (Guo et al., 2022), and CodeT5+ embedding (Wang et al., 2023). In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3  (Sturua et al., 2024), All-MiniLM-L12-v2 (Reimers and Gurevych, 2021), All-mpnet-base-v2  (Song et al., 2021), and Multilingual-e5-large  (mul, 2023).
@@
Method NDCG@10 MRR
BoW 0.046 0.034
CodeBERT 0.232 0.175
UniXcoder 0.187 0.140
CodeT5+ embedding 0.179 0.136
Jina-embedding-v3 0.225 0.171
All-MiniLM-L12-v2 0.214 0.163
All-mpnet-base-v2 0.227 0.172
Multilingual-e5-large 0.218 0.166
Table 4: Performance of Various Methods on CoSQA+ Using NDCG@10 and MRR Metrics.
While traditional information retrieval methods, such as BoW, are computationally efficient, Table 4 demonstrates that they struggle to capture the semantic subtleties needed to accurately map natural language queries to relevant code snippets, particularly when multiple valid solutions exist. Among the deep learning-based models, CodeBERT achieves the highest performance with an NDCG@10 of 0.232 and an MRR of 0.175. This strong performance can be attributed to its encoder-only architecture and extensive pre-training on large-scale paired natural and programming language corpora, especially Python, which enhances its code understanding capabilities. Other embedding-based models also show substantial improvements over traditional methods, highlighting the benefits of modern deep learning architectures for semantic code search.
@@
Conclusion
In this paper, we present CoSQA+, the first large-scale multi-choice code search dataset with 412,080 pairs. Our dataset expands from single-choice to multi-choice query-code matching pairs, enhancing its suitability for real-world code search tasks. Through rigorous experiments, CoSQA+ demonstrates superior quality and training performance compared to existing benchmarks like CoSQA. The test-driven annotation approach achieves a high accuracy of 96.4%, outperforming traditional human annotation and single LLM methods. Experiments on CoSQA+ show that CodeBERT is better suited for multi-choice code search tasks. We believe our agents-annotated CoSQA+ benchmark will offer insights for other research that involves natural language and code, and for evaluations on a multi-choice code search task
Ethical consideration

### v5

@@ IV-B Candidate Pairs Construction
The process of constructing candidate codes for queries in CoSQA+ involves 3 steps aimed at efficiently filtering potential matching codes from a large codebase. This forms the foundation for annotating query-code pairs in the CoSQA+ dataset. Firstly, we employ three well-established code embedding models to generate vector representations for both queries and codes. Then, these embeddings are used to calculate cosine similarity scores, averaging similarity results across the three models. Finally, the top 20 codes with the highest similarity scores are selected as candidates for each query. Fig.  3 shows a detailed breakdown of query distributions by the number of exactly matched codes.
Figure 3: Query counts grouped by number of matched codes
@@
High similarity between query and code is fundamental for a high-quality dataset [57]. During our preliminary assessments, we observe disparities in the similarity scores calculated by different models. Relying solely on one model for computing similarity could introduce bias and skew the dataset unfavorably. This potentially leads to unfair test results of the benchmark. To mitigate this bias, we employ a multi-model approach for calculating the similarity of query-code pairs. We empirically validate this practice. Specifically, we use a single model to select 20 candidate codes for each query and use an ensemble of models, CodeBERT [12], UniXcoder [31] and CodeT5+ 110M embedding [34], to select another 20 candidates. Each model embeds each code and query as a vector based on different underlying mechanisms and training datasets. We calculate the overlap between the candidates chosen by the single model and those chosen by the ensemble and analyze the overlap between them. Although bias cannot be completely eliminated, the experiment result in Table I shows that comprehensively utilizing multiple models for matching can effectively reduce the bias of query-code pairs toward any single model. As can be seen, the proportion of overlapping codes is low, indicating that the codes matched by multiple models for a query do not show a significant preference for any particular single model.
TABLE I: Overlap ratio of candidate codes between individual models and the ensemble of models.
Model Overlap Ratio
CodeBERT 0.135
UniXcoder 0.164
CodeT5+ embedding 0.088
@@
By taking the average similarity score from these three models, we achieve a more balanced and fair assessment of query-code pair quality.
@@
Before constructing the full set of candidate pairs, we randomly select 2,000 queries and annotated them using test-driven agents to determine the optimal number of candidates per query. The experimental results indicate that the probability of identifying relevant codes decreases significantly as the number of candidates increases (Fig. 4). As shown in our analysis, when the number of candidates reaches 20, the marginal benefit of further expansion diminishes. This decline can be attributed to limitations of codebase, nature of queries and ability of embedding models. Furthermore, adding just one more candidate code per query results in an additional 20,604 pairs, which would incur substantial costs in terms of both time and financial resources for annotation. Thus, limiting candidates to the top-20 strikes a balance between maximizing the retrieval of relevant codes and controlling annotation costs.
@@
IV-C Test-Driven Agents Annotation
To efficiently and accurately annotate the large number of query-code pairs, we design the test-driven annotation agents based on DeepSeek-V3 [60], using the algorithm shown in Algorithm 1. Initially, A preliminary screener determines whether the query clearly matches or does not match the code, or if it is unclear For clear cases, we directly assign labels based on the initial assessment. For uncertain cases, we determine the labels based on the test execution results and final judgment. In both paths, labels are assigned to indicate whether the code matches the query’s requirements (1) or not (0). All labeled query-code pairs are collected to form our dataset CoSQA+, which consists of pairs that have been rigorously verified either through direct assessment or test-based validation.
Algorithm 1 Test-Driven Agent Annotation
Input: query $q$, code snippet $c$, max_try_time
Output: label $l\in\{0,1\}$
1:  $assessment\leftarrow\text{preliminary\_screener}(q,c)$
[…]
3:     return  1
[…]
5:     return  0
[…]
7:     $test\_prog\leftarrow\text{generate\_test\_program}(q,c)$
8:     $exec\_res\leftarrow\text{execute\_test}(test\_prog)$
9:     $try\_time\leftarrow 0$
10:     while $exec\_res.\text{error}\ \textbf{and}{}\ try\_time\leq max\_try\_time$ do
11:        $terminal\_commands\leftarrow\text{fix\_test\_program}(test\_prog,exec\_res.\text{error})$
[…]
13:        $exec\_res\leftarrow\text{execute\_test}(test\_prog)$
14:        $try\_time\leftarrow try\_time+1$
15:     end while
[…]
17:  end if
@@ IV-C Test-Driven Agents Annotation
The evaluation process begins with a preliminary screener that categorizes query-code pairs into three distinct labels: (1) clear match, (2) clear non-match, and (3) unclear, as illustrated in Fig. 5. Guided by the prompt instructions, the model assesses whether the code fully addresses the query’s functional requirements. For unambiguous cases, it provides a justification and immediate classification. For ambiguous cases (representing 33% of candidates), it flags the pair for deeper verification by downstream agents. This triage mechanism optimizes the annotation workflow by resolving straightforward cases upfront while reserving computational resources for ambiguous scenarios.
@@
For ambiguous cases, we employ a test-driven verification methodology. The test program generator constructs executable test suites designed to rigorously validate code functionality. Each test suite includes dependency scaffolding — importing required libraries, defining auxiliary classes, and declaring helper functions — to ensure runtime compatibility. The tests incorporate assert-based validation mechanisms targeting edge cases, functional requirements, and behavioral constraints specified in the query. Considering that a single query may have multiple valid explanations, we generate the test program based on one of the plausible interpretations of the query. This approach transforms ambiguous cases into executable specifications for objective validation.
@@
The framework ultimately concludes with a final arbiter stage, where the agent employs a specialized prompt to assess if the generated code fully satisfies the query. This assessment is based on both the test program produced and the outcomes of its execution. This methodology shifts verification from theoretical code inspection to empirical observation, empowering the agent to establish evidence-based conclusions by systematically evaluating actual runtime behavior rather than relying exclusively on static structural analysis and semantic analysis. Fig. 6 illustrates the process of handling an unclear case.
Figure 6: An example for handling an unclear case through test program generation, test execution, and final arbitration.
IV-D Benchmark Statistics
In total, 412,080 labeled query-code pairs are retained. Table II shows the statistics of our CoSQA+ benchmark.
TABLE II: Statistics of CoSQA+. "#" denotes the count, "Code avg. len." is the average number of characters per code snippet, and "Exec. rate" represents the percentage of test programs that executed successfully.
# of pairs # of code Code avg. len. Avg. test cases Exec. rate
@@ IV-D Benchmark Statistics
IV-E Cost Analysis
@@
In contrast, manual annotation by Python experts involves substantial time investment. On average, a Python expert spends 8 minutes per sample due to extensive research and analysis. For test-driven annotation, where experts design tailored test programs for code-query pairs, the time increases to 34 minutes per sample, with complex functions exceeding one hour.
IV-F Informed consent
This study involved human participants. Informed consent was obtained from all participants prior to their involvement in the study. Participants were informed about the purpose of the research, the procedures involved, the potential risks and benefits, and their right to withdraw from the study at any time without penalty. All data collected were anonymized to ensure participant confidentiality.
V Evaluation
@@
RQ4: How does each agent contribute to the performance of test-driven agents?
V-A RQ1: Quality Comparison between CoSQA+ and CoSQA
@@
Following previous work, we employ two metrics to evaluate the performance of code search models: NDCG [14] and MRR [12]. NDCG measures the quality of ranking by considering both the relevance and position of retrieved results. For a given query $q_{i}$, we calculate NDCG@k as:
@@
In the first experiment, we randomly select 1000 query-code pairs labeled "1" from CoSQA+ and CoSQA, ensuring the queries are identical. For CoSQA+, a query may have multiple correct codes, from which we randomly select one that differs from the corresponding query’s code in CoSQA. These queries are then presented to experienced programmer volunteers, who are asked to choose the code that satisfies more demand of the query. Finally, we calculate the proportion of better codes from CoSQA+ and CoSQA.
@@
TABLE III: Performance of code search models fine-tuned with different data. The test set is CSN Python.
@@
After the volunteers compare the two datasets, we come to the result that out of the 2000 selected queries, 60.3% of the better codes are from CoSQA+, while 39.7% are from CoSQA. This indicates that our CoSQA+ is superior in terms of code quality and the degree of matching between queries and codes. As shown in Table III, the three models fine-tuned with CoSQA+ outperform those fine-tuned with CoSQA, which indicates better training performance of CoSQA+. Notably, the improvement in NDCG@10 is more significant than the improvement in MRR, highlighting the effectiveness of CoSQA+ in training the model to retrieve multiple relevant code matches for a single query.
@@
V-B RQ2: Performance of Agent-based Annotator
@@
To establish a reliable ground truth, we first have three Python experts design test programs for 400 randomly selected query-code pairs and annotate them based on test execution results. The final ground truth labels are determined by majority voting among these three experts. We then evaluate three different annotation approaches against this ground truth: (1) human annotators without test program verification; (2) our agent-based annotator; (3) single LLM annotators (including Deepseek-V3, O1-mini, and Claude-3.5-Sonnet).
For measuring the agreement across three Python experts with test programs, we use Krippendorff’s alpha coefficient, which is particularly suitable for assessing agreement among multiple annotators [61].
@@
TABLE IV: Accuracy comparison of different annotation methods against ground truth.
@@
Agent-based Annotator 0.920
Human Wo. Test Program 0.880
Deepseek-V3 0.865
O1-mini 0.833
Claude-3.5-Sonnet 0.845
As shown in Table IV, our agent-based annotator achieves 92.0% accuracy compared to the ground truth, significantly outperforming human annotators without test program verification (89.0%), Deepseek-V3 (84.0%), O1-mini (82.0%) and Claude-3.5-Sonnet (84.0%). This high accuracy demonstrates that our automated agent-based approach can effectively match the quality of expert annotations with test program verification.
@@
In terms of inter-annotator agreement measured by Krippendorff’s alpha coefficient, the experts with test programs show strong consistency with a coefficient of 0.818, notably higher than human annotations without test program verification (0.61), which suggests that test programs help reduce subjectivity in the annotation process.
RQ2 Summary: Our agent-based annotator achieves 92.0% accuracy, outperforming both traditional human annotation and single LLM approaches. The high executability rate (83.67%) of generated test programs demonstrates the reliability of our automated annotation approach.
V-C RQ3: Existing Methods Performance on CoSQA+
@@
The effectiveness of semantic code search models is evaluated on the CoSQA+ benchmark using NDCG@10 as the primary metric. NDCG@10 is particularly suitable for our scenario, where multiple correct code snippets may exist for a query, as it measures how well a model ranks all relevant results rather than focusing only on the first correct match. Additionally, we report Mean Reciprocal Rank (MRR) to assess the ability to rank the first relevant result highly. To further comprehensively evaluate the performance of the models, we also introduce Mean Average Precision (MAP) and Recall as additional metrics. MAP provides a measure of the overall ranking quality by averaging the precision scores across all relevant results, which is especially useful for scenarios where the order of multiple correct matches matters. Meanwhile, Recall evaluates the model’s ability to retrieve all relevant code snippets from the dataset, ensuring that no important results are missed. Together, these metrics offer a more robust assessment of the model’s effectiveness in semantic code search tasks. The tested methods include both traditional information retrieval techniques and advanced deep learning-based models. Specifically, we compare a traditional Bag of Words (BoW) approach [27] with state-of-the-art semantic code search models, namely: CodeBERT [12], UniXcoder [31], and CodeT5+ embedding [34]. In addition, we evaluate several recent sentence embedding models, including Jina-embedding-v3 [35], All-MiniLM-L12-v2 [40], All-mpnet-base-v2 [41], and Multilingual-e5-large [38].
@@
TABLE V: Performance of Various Methods on CoSQA+ Using NDCG@10, MRR, MAP and Recall Metrics.
Method NDCG@10 MRR MAP Recall
BoW 0.037 0.027 0.027 0.016
CodeBERT 0.232 0.175 0.164 0.149
UniXcoder 0.187 0.140 0.132 0.110
CodeT5+ embedding 0.179 0.136 0.127 0.101
Jina-embedding-v3 0.225 0.171 0.160 0.145
All-MiniLM-L12-v2 0.214 0.163 0.153 0.135
All-mpnet-base-v2 0.227 0.172 0.162 0.145
Multilingual-e5-large 0.218 0.166 0.156 0.136
While traditional information retrieval methods, such as BoW, are computationally efficient, Table V demonstrates that they struggle to capture the semantic subtleties needed to accurately map natural language queries to relevant code snippets, particularly when multiple valid solutions exist. Among the deep learning-based models, CodeBERT achieves the highest performance with NDCG@10 of 0.232, MRR of 0.175, MAP of 0.164 and Recall of 0.149. Other embedding-based models also show substantial improvements over traditional methods, highlighting the benefits of modern deep learning architectures for semantic code search. This strong performance can be attributed to their encoder-only architecture and extensive pre-training on large-scale paired natural and programming language corpora, especially Python, which enhances their code understanding capabilities. The embedding models originate from different time periods and vary in parameter sizes. Surprisingly, CodeBERT outperforms the latest embedding models, which may be attributed to the fact that CodeBERT tested here is fine-tuned specifically for code search tasks, whereas the latest embedding models are designed for general-purpose tasks. Some smaller-parameter embedding models demonstrate impressive performance. For instance, all-mpnet-base-v2, with only 109M parameters, and all-MiniLM-L12-v2, with a mere 33M parameters, still exhibit remarkable performance, surpassing many models with larger parameter sizes. This represents a significant advancement in embedding models in recent years, providing a solid foundation for the practical implementation of semantic search based on large models.
From the evaluation results, the performance differences among embedding models seem not significant. However, due to the richness of CoSQA+, which effectively eliminates random factors, even subtle differences in metrics can reflect the overall capability gap between models. Moreover, embedding models generally perform poorly on our benchmark. From the perspective of model design, embedding models are mainly used for retrieval tasks. To improve inference speed, their weight parameters are usually small, limiting their ability to deeply understand query sentences and code.
The CoSQA+ is challenging in two main aspects: First, the queries originate from CoSQA, which are real user queries filtered from the Bing search engine. These queries are often expressed casually and may contain spelling or grammar errors. Second, the codebase used for retrieval is not only large in scale but also contains many confusing code snippets.
@@
V-D RQ4: Contributions of Different Components in Test-Driven Agents
This experiment is to assess the contribution of each component in our test-driven agent framework.
Experimental setup
The final test-driven agents comprise the following components: Preliminary Screener (PS), Test Program Generator (TPG), Test Executor (TE), Bug Fixer (BF), and Final Arbiter (FA). We divided the agent into different configurations by combining these components in various ways. For each configuration, we compared the predicted labels with the ground truth labels (obtained from human evaluation based on the test programs) to calculate the accuracy. Additionally, we compare the Simulated Executor with both the Test Executor and the Bug Fixer to evaluate whether the large language model can accurately simulate execution without actual interpretation. Unlike the Test Executor, which runs the test program using a Python interpreter, the Simulated Executor leverages the LLM’s reasoning capabilities to emulate execution and directly predict the outcome.
Result and Analysis
Table VI shows the accuracy for different configurations.
Configuration Accuracy
Single LLM 0.865
PS + FA 0.840
PS + TPG + FA 0.833
PS + TPG + TE + FA 0.835
PS + TPG + Simulated Executor + FA 0.840
PS + TPG + TE + BF + FA 0.920
TABLE VI: Accuracy of Various Test-Driven Agent Configurations (Abbreviated)
The results indicate that the baseline Single LLM achieves an accuracy of 0.865. When the system uses only the Preliminary Screener in combination with the Final Arbiter, the accuracy slightly drops to 0.840. Incorporating the Test Program Generator alongside these yields an accuracy of 0.833, suggesting that generating test programs without execution may introduce noise or irrelevant information, which can harm performance [62]. Adding the Test Executor to the configuration improves the performance marginally (0.835), and using a Simulated Test Executor does not provide a significant advantage (0.840). This aligns with the observation that simulated execution cannot fully capture the complexity of the real compile-and-run process, potentially introducing inaccuracies or misleading feedback. Notably, the inclusion of the Bug Fixer increases the accuracy substantially to 0.920. This enhancement demonstrates that the Bug Fixer is crucial for processing the test programs effectively, as it enables the system to generate real execution results that are vital for improving prediction accuracy.
RQ4 Summary: The Bug Fixer is the most important component in the test-driven agent framework as it makes the test program executable and yields effective test results. However, it relies on the preceding steps, such as generating the test program. Simulated execution results cannot replace the benefits of real execution feedback.
VI Discussion
VI-A Test Misjudgment
Previous studies show that as the size of the codebase increases, the challenges faced by embedding models also increase [63, 64]. This conclusion aligns with intuition, but with the analysis of other datasets and the construction of CoSQA+, we find that the decline in performance is not only due to the need to exclude more confusing codes but also because of the increased likelihood of test misjudgments. For example, a LLM may select a piece of code that meets the query requirements, but this query-code pair is not included in the candidate code set. As a result, this correct selection is misjudged as incorrect.
When constructing code search datasets, query sets and code sets are typically extracted from original query-code pairs for test. However, the number of all possible query-code pairs (similar to a Cartesian product) formed between the query set and the code set far exceeds the number of original pairs. Existing methods assume that the code of a pair outside the original pairs cannot correctly match the query, but this assumption is not entirely valid. As shown in Fig. 7, in CoSQA+, excluding codes without satisfied query, 36.7% of codes can satisfy two or more queries. Constructing a dataset completely free of misjudgments is almost impossible, as it would require checking all possible pairs. As can be seen by combining Fig 3 and Fig. 7, the approach adopted in CoSQA+, where 20 candidate codes are annotated for each query, helps mitigate these testing misjudgments. Another possible measure is to invoke test-driven annotation agents when a code search method selects a previously unseen query-code pair to determine whether the selected code matches the query. However, this approach may introduce significant computational overhead, especially when the candidate code repository is very large.
Figure 7: Distribution of codes by the number of queries they satisfy
VI-B Inherent Ambiguity in Query and Code Labels
A major challenge in code search is the inherent ambiguity in queries and code. Queries are often vague or poorly phrased, leading to multiple interpretations, while code snippets may be incomplete or abstracted, making correctness subjective. For instance, a query like "how to check if a folder exists and create it if not" can be interpreted differently based on assumptions about paths or edge cases. Similarly, code may solve the problem partially or in a non-standard way, complicating strict labeling ("exact match"). Test-driven annotation helps reduce ambiguity by generating executable test programs aligned with the query’s intent. However, variability in query interpretation and incomplete code still introduce labeling challenges. Future improvements will require clearer queries and more complete code snippets to minimize ambiguity.
VI-C Ethical Consideration